A force of magnitude 5unit in the direction of vector i^+j^+k^ is displaindi the particle from point (1,1,1) to (2,1,3). Find work done by force.
 

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
as work done = F . S
 
F = 5 (i + j+ k) / sqrt(3)
 
Also displacement vector = i + 2k
 
hence work done = F. S = 5/ sqrt(3) + 10 / sqrt(3) = 15 / sqrt(3) = 5* sqrt(3)